OpenStudy (anonymous):

differentiate:xy'+2y=x^2

OpenStudy (anonymous):

I have one solution: c/x^2+x^2/4; but confusion, could any one help me

OpenStudy (amistre64):

define for me "differentiate" as it applies to this problem please...

OpenStudy (amistre64):

...ordinary differential equation....right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

if I recall correctly, that means your trying to find the original function that this was derived from correct?

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

the only method I remember off hand is the seperation of variables.... have you tried that yet?

OpenStudy (anonymous):

i think seperation variable is not correct for this equation

OpenStudy (amistre64):

youre probably right...step me through what youve done already

OpenStudy (anonymous):

i apply bernoulli's equa.

OpenStudy (amistre64):

y' + P(x)y = Q(x)y^n.. that one?

OpenStudy (anonymous):

y'+p(x)y=r(x)

OpenStudy (amistre64):

hmmm.... I havent had much practice with ode's .... maybe someone smarter will come along :)

OpenStudy (anonymous):

You can use the method: multiplying with an integrating factor. divide everything by x, so you'll have y'+2/x y=x. The integrating factor will then be: e^{2\int 1/x}=e^{2lnx}=e^{ln(x^2)}=x^2.

OpenStudy (anonymous):

Multiply both sides by the integrating factor x^2, then the left side will be (y*x^2)'. The right side is x^3. Then you integrate both sides, and gets y*x^2={1/4} *x^4+C Divide both sides of the equation by x^2 and you have the answer :) If you're not familiar with the method and wants a further explanation, just tell me...

OpenStudy (anonymous):

c/x^2+x^2/4 i have already mention above

OpenStudy (anonymous):

is it correct?

OpenStudy (anonymous):

Yes, it is correct. y=... , youre right. I only looked too much at the latter posts instead of where your answer was, I'm sorry for that..